OpenStudy (anonymous):

what is the simplified form of x/3-y/4/x/4+y/3

zepdrix (zepdrix):

Is this what the problem looks like?\[\large \frac{x}{3}-\frac{\left(\dfrac{y}{4}\right)}{\left(\dfrac{x}{4}\right)}+\frac{y}{3}\] The brackets are to show which fraction belongs where, otherwise the problem doesn't make sense D:

OpenStudy (anonymous):

yessir @zepdrix

zepdrix (zepdrix):

When dividing fractions, we can rewrite it as multiplication by flipping the bottom fraction.\[\large \frac{\left(\dfrac{y}{4}\right)}{\left(\dfrac{x}{4}\right)} \qquad=\qquad \frac{y}{4}\cdot \frac{4}{x}\]

zepdrix (zepdrix):

The 4's will divide out giving us,\[\large \frac{4y}{4x}=\frac{y}{x}\] Simplifying the middle term leaves us with,\[\large \frac x3-\frac yx+\frac y3\]

zepdrix (zepdrix):

From here, you need to get a common denominator. In this case it will be 3 times x. \[\large \left(\frac{x}{x}\right)\frac x3-\left(\frac{3}{3}\right)\frac yx+\left(\frac{x}{x}\right)\frac y3\] \[\large \frac{x^2}{3x}-\frac{3y}{3x}+\frac{xy}{3x} \qquad=\qquad \frac{x^2+xy-3y}{3x}\] Hmm this is a weird problem :O you sure we had the fractions correct at the start?

OpenStudy (anonymous):

yeaaaap lol